# Function theory on the annulus in the dp-norm

### v1

@@ 5. The DP Pick problem and DP-Szegő kernels
Suppose the kernel $[\frac{1}{1-\bar{\mu}\lambda}]$ restricted to $R_{\delta}$ is a DP kernel. Then, for any distinct $\lambda_{1},\dots,\lambda_{n}\in R_{\delta}$, the localization of $[\frac{1}{1-\bar{\mu}\lambda}]$ to $\{\lambda_{1},\dots,\lambda_{n}\}$ belongs to $\mathcal{G}_{\mathrm{dp}}(\lambda)$.
Consider the $2$ distinct points $\lambda_{1}=\frac{\delta+\sqrt{\delta}}{2}$, $\lambda_{2}=-\lambda_{1}$, note that, for $\delta\in(0,1)$, $\lambda_{1},\lambda_{2}\in R_{\delta}$. By Example (2.9), the function $\varphi(z)=\tfrac{1}{2}(z+\frac{\delta}{z})$ on $R_{\delta}$ satisfies $\|\varphi\|_{\mathrm{dp}}=1$. Therefore, for $\lambda_{i}$ and $z_{i}=\varphi(\lambda_{i})=\frac{1}{2}\left(\lambda_{i}+\frac{\delta}{\lambda_{i}}\right)$, $i=1,2$, the DP Pick problem
@@
$[(1-\bar{z}_{i}z_{j})g_{ij}]\geq 0.$ (5.26)
By the assumption, the localization of $[\frac{1}{1-\bar{\mu}\lambda}]$ to $\{\lambda_{1},\lambda_{2}\}$ belongs to $\mathcal{G}_{\mathrm{dp}}(\lambda)$. In particular, for the Pick problem
@@
$\left[\frac{1-\bar{z}_{i}z_{j}}{1-\bar{\lambda}_{i}\lambda_{j}}\right]_{i,j=1}^{2}\geq 0.$ (5.28)
Hence, by Pick’s theorem, the problem (5.27) is solvable by a Schur function $f$ on $\mathbb{D}$. This contradicts our example, and so $[\frac{1}{1-\bar{\mu}\lambda}]$ is not a DP Szegő kernel on $R_{\delta}$. ∎
@@ 6. Extremal DP Pick problems
We say that a DP Pick problem is extremally solvable if it is solvable and there does not exist $\varphi\in{\rm H}^{\infty}_{\rm dp}$ with $\|\varphi\|_{\rm dp}<1$ satisfying the equations
@@
    $[(1-\bar{z}_{i}z_{j})g_{ij}]_{i,j=1}^{n}\geq 0$ (6.4)
    and there exists $\tilde{g}\in\mathcal{G}_{\rm dp}(\lambda)$ such that
    $\operatorname{rank}\big{[}(1-\bar{z}_{i}z_{j})\tilde{g}_{ij}\big{]}_{i,j=1}^{n}<n.$ (6.5)
@@
[…]
$=\min_{J=1,\dots,n}\det\begin{bmatrix}(1-r^{2}\bar{z}_{i}z_{j})g_{ij}\end{bmatrix}_{i,j=1}^{J}.$
By standard linear algebra, for any positive definite matrix $g=[g_{ij}]$, $F(r,g)>0$ if and only if $\begin{bmatrix}(1-r^{2}\bar{z}_{i}z_{j})g_{ij}\end{bmatrix}_{i,j=1}^{n}>0$. $F$ is continuous and, by supposition,
$[(1-\bar{z}_{i}z_{j})g_{ij}]>0$
@@
By the continuity of $F$ and, again by the compactness of $\mathcal{G}_{\mathrm{dp}}^{\mathrm{norm}}(\lambda)$, the family of functions $\{F(\cdot,g):g\in\mathcal{G}_{\mathrm{dp}}^{\mathrm{norm}}(\lambda)\}$ is equicontinuous on $\mathbb{R}$. Hence there exists $\delta>0$ such that $F(r,g)>0$ for all $g\in\mathcal{G}_{\mathrm{dp}}^{\mathrm{norm}}(\lambda)$ and all $r\in(1,\delta)$. Choose some $r\in(1,\delta)$. Then $F(r,g)>0$ for all $g\in\mathcal{G}_{\mathrm{dp}}^{\mathrm{norm}}(\lambda)$, and therefore $\begin{bmatrix}(1-r^{2}\bar{z}_{i}z_{j})g_{ij}\end{bmatrix}_{i,j=1}^{n}>0$ for all $g\in\mathcal{G}_{\mathrm{dp}}^{\mathrm{norm}}(\lambda)$. It follows from Proposition 4.24 that $\begin{bmatrix}(1-r^{2}\bar{z}_{i}z_{j})g_{ij}\end{bmatrix}_{i,j=1}^{n}>0$ for all $g\in\mathcal{G}_{\mathrm{dp}}(\lambda)$. Hence, by Theorem 5.2, for the chosen $r\in(1,\delta)$, the DP Pick problem
@@
$[(r^{2}-\bar{z_{i}}z_{j})g_{ij}]\geq 0.$ (6.11)
By assumption (ii), there exists $\tilde{g}\in\mathcal{G}_{\rm dp}(\lambda)$ such that
$\operatorname{rank}\big{[}(1-\bar{z}_{i}z_{j})\tilde{g}_{ij}\big{]}_{i,j=1}^{n}<n.$ (6.12)
and hence $[(1-\bar{z}_{i}z_{j})\tilde{g}_{ij}]_{i,j=1}^{n}$ has a non-zero null vector $v$. Consider the relation
$(1-\bar{z}_{i}z_{j})\tilde{g}_{ij}=(1-r^{2})\tilde{g}_{ij}+(r^{2}-\bar{z}_{i}z_{j})\tilde{g}_{ij}.$
Since $g_{ij}>0$, $(1-r^{2})\tilde{g}_{ij}>0$ and, by equation (6.11), $(r^{2}-\bar{z}_{i}z_{j})\tilde{g}_{ij}\geq 0$, which is a contradiction. ∎
@@
Let us now prove the existence of an operator $T$ with the stated properties. By assumption, the DP-Pick problem $\lambda_{j}\mapsto z_{j},j=1,\dots,n$ is extremally solvable. By Theorem 6.3, there exists $\tilde{g}\in\mathcal{G}_{\rm dp}(\lambda)$ such that
$\operatorname{rank}\big{[}(1-\bar{z}_{i}z_{j})\tilde{g}_{ij}\big{]}_{i,j=1}^{n}<n,$ (6.16)
so that $\begin{bmatrix}(1-\bar{z}_{i}z_{j})\tilde{g}_{ij}\end{bmatrix}$ is singular, and therefore has a non-zero null vector $\xi=[\xi_{1},\dots,\xi_{n}]^{T}\in\mathbb{C}^{n}$, which is to say that
$\sum_{j=1}^{n}(1-\bar{z}_{i}z_{j})\tilde{g}_{ij}\xi_{j}=0\ \text{for}\ i=1,\dots,n.$ (6.17)
Since $\tilde{g}\in\mathcal{G}_{\rm dp}(\lambda)$, $[\tilde{g}_{ij}]>0$, and so $[\tilde{g}_{ij}]$ has rank $n$. By Moore’s Theorem there exist an $n$-dimensional Hilbert space $\mathcal{H}$ and a basis $\tilde{e}_{1},\dots,\tilde{e}_{n}\in\mathcal{H}$ such that $\tilde{g}_{ij}=\langle\tilde{e}_{j},\tilde{e}_{i}\rangle$ for $i,j=1,\dots,n$.
Define an operator $T$ on $\mathcal{H}_{T}$ by $T\tilde{e}_{j}=\lambda_{j}\tilde{e}_{j}$ for $j=1,\dots,n$. Since $\tilde{g}\in\mathcal{G}_{\rm dp}(\lambda)$, by Proposition 4.10, $T\in\mathcal{F}_{\rm dp}(\delta,\lambda)$. Note that $\varphi(T)\tilde{e}_{j}=z_{j}\tilde{e}_{j}$, and so, if $x=\sum_{j=1}^{n}\xi_{j}\tilde{e}_{j}$, then
$\varphi(T)x=\sum_{j=1}^{n}z_{j}\xi_{j}\tilde{e}_{j}$
@@
$\langle(1-\varphi(T)^{*}\varphi(T))x,x\rangle$ $=\sum_{i,j=1}^{n}(1-\bar{z}_{i}z_{j})\bar{\xi}_{i}\xi_{j}\langle\tilde{e}_{j},\tilde{e}_{i}\rangle$
$=\sum_{i,j=1}^{n}(1-\bar{z}_{i}z_{j})\bar{\xi}_{i}\xi_{j}\tilde{g}_{ij}$
$=\sum_{i=1}^{n}\bar{\xi}_{i}\sum_{j=1}^{n}(1-\bar{z}_{i}z_{j})\xi_{j}\tilde{g}_{ij}$
@@
As $\xi\neq 0$, the complex numbers $\xi_{1},\dots,\xi_{n}$ are not all zero, and so, since $\tilde{e}_{1},\dots,\tilde{e}_{n}$ are linearly independent, $x=\sum_{j=1}^{n}\xi_{j}\tilde{e}_{j}\neq 0$. Since $\|\varphi\|_{\mathrm{dp}}\leq 1$ and $T\in\mathcal{F}_{\rm dp}(\delta,\lambda)$, we have $\|\varphi(T)\|\leq 1$, and so $1-\varphi(T)^{*}\varphi(T)\geq 0$. In conjunction with the inequality (6), this implies that $(1-\varphi(T)^{*}\varphi(T))x=0$, and hence $\|\varphi(T)x\|^{2}=\|x\|^{2}$. Since $x\neq 0$, $x$ is a maximizing vector for $\varphi(T)$ and $\|\varphi(T)\|=1$.
@@
where $\dim\mathcal{M}\leq 2n$. Since $T\in\mathcal{F}_{\rm dp}(\delta,\lambda)$, $T$ satisfies
@@
$1-\varphi(T)^{*}\varphi(T)=u(T)^{*}\big{(}1-E(T)^{*}E(T)\big{)}\ u(T).$

### v2

@@ 5. The DP Pick problem and DP-Szegő kernels
Suppose the kernel $\frac{1}{1-\overline{\mu}\lambda}]$ restricted to $R_{\delta}$ is a DP kernel. Then, for any distinct $\lambda_{1},\dots,\lambda_{n}\in R_{\delta}$, the localization of $\frac{1}{1-\overline{\mu}\lambda}]$ to $\{\lambda_{1},\dots,\lambda_{n}\}$ belongs to $\mathcal{G}_{\mathrm{dp}}(\lambda)$.
Consider the $2$ distinct points $\lambda_{1}=\frac{\delta+\sqrt{\delta}}{2}$, $\lambda_{2}=-\lambda_{1}$, note that, for $\delta\in(0,1)$, $\lambda_{1},\lambda_{2}\in R_{\delta}$. By Example 2.9, the function $\varphi(z)=\tfrac{1}{2}(z+\frac{\delta}{z})$ on $R_{\delta}$ satisfies $\|\varphi\|_{\mathrm{dp}}=1$. Therefore, for $\lambda_{i}$ and $z_{i}=\varphi(\lambda_{i})=\frac{1}{2}\left(\lambda_{i}+\frac{\delta}{\lambda_{i}}\right)$, $i=1,2$, the DP Pick problem
@@
$(1-\overline{z}_{i}z_{j})g_{ij}]\geq 0.$ (5.29)
By the assumption, the localization of $\frac{1}{1-\overline{\mu}\lambda}]$ to $\{\lambda_{1},\lambda_{2}\}$ belongs to $\mathcal{G}_{\mathrm{dp}}(\lambda)$. In particular, for the Pick problem
@@
$\left[\frac{1-\overline{z}_{i}z_{j}}{1-\overline{\lambda}_{i}\lambda_{j}}\right]_{i,j=1}^{2}\geq 0.$ (5.31)
Hence, by Pick’s theorem, the problem (5.30) is solvable by a Schur function $f$ on $\mathbb{D}$. This contradicts our example, and so $\frac{1}{1-\overline{\mu}\lambda}]$ is not a DP Szegő kernel on $R_{\delta}$. ∎
@@ 6. Extremal DP Pick problems
In this section we study DP Pick interpolation problems that are “only just” solvable. We say that a DP Pick problem is extremally solvable if it is solvable and there does not exist $\varphi\in{\rm H}^{\infty}_{\rm dp}$ with $\|\varphi\|_{\rm dp}<1$ satisfying the equations
@@
Next we give necessary and sufficient conditions for a DP Pick problem to be extremally solvable.
@@
    $(1-\overline{z}_{i}z_{j})g_{ij}]_{i,j=1}^{n}\geq 0$ (6.4)
    and there exists $\widetilde{g}\in\mathcal{G}_{\rm dp}(\lambda)$ such that
    $\operatorname{rank}\big[(1-\overline{z}_{i}z_{j})\widetilde{g}_{ij}\big]_{i,j=1}^{n}<n.$ (6.5)
@@
$F(r,[g_{ij}])$ $=\text{the minimum of the leading principal minors of}\begin{bmatrix}(1-r^{2}\overline{z}_{i}z_{j})g_{ij}\end{bmatrix}_{i,j=1}^{n}$
$=\min_{J=1,\dots,n}\det\begin{bmatrix}(1-r^{2}\overline{z}_{i}z_{j})g_{ij}\end{bmatrix}_{i,j=1}^{J}.$
By standard linear algebra, for any positive definite matrix $g=[g_{ij}]$, $F(r,g)>0$ if and only if $\begin{bmatrix}(1-r^{2}\overline{z}_{i}z_{j})g_{ij}\end{bmatrix}_{i,j=1}^{n}>0$. $F$ is continuous and, by supposition,
$(1-\overline{z}_{i}z_{j})g_{ij}]>0$
@@
By the continuity of $F$ and, again by the compactness of $\mathcal{G}_{\mathrm{dp}}^{\mathrm{norm}}(\lambda)$, the family of functions $\{F(\cdot,g):g\in\mathcal{G}_{\mathrm{dp}}^{\mathrm{norm}}(\lambda)\}$ is equicontinuous on $\mathbb{R}$. Hence there exists $\delta>0$ such that $F(r,g)>0$ for all $g\in\mathcal{G}_{\mathrm{dp}}^{\mathrm{norm}}(\lambda)$ and all $r\in(1,\delta)$. Choose some $r\in(1,\delta)$. Then $F(r,g)>0$ for all $g\in\mathcal{G}_{\mathrm{dp}}^{\mathrm{norm}}(\lambda)$, and therefore $\begin{bmatrix}(1-r^{2}\overline{z}_{i}z_{j})g_{ij}\end{bmatrix}_{i,j=1}^{n}>0$ for all $g\in\mathcal{G}_{\mathrm{dp}}^{\mathrm{norm}}(\lambda)$. It follows from Proposition 4.23 that $\begin{bmatrix}(1-r^{2}\overline{z}_{i}z_{j})g_{ij}\end{bmatrix}_{i,j=1}^{n}>0$ for all $g\in\mathcal{G}_{\mathrm{dp}}(\lambda)$. Hence, by Theorem 5.2, for the chosen $r\in(1,\delta)$, the DP Pick problem
@@
$(r^{2}-\overline{z_{i}}z_{j})g_{ij}]\geq 0.$ (6.11)
By assumption (ii), there exists $\widetilde{g}\in\mathcal{G}_{\rm dp}(\lambda)$ such that
$\operatorname{rank}\big[(1-\overline{z}_{i}z_{j})\widetilde{g}_{ij}\big]_{i,j=1}^{n}<n.$ (6.12)
and hence $(1-\overline{z}_{i}z_{j})\widetilde{g}_{ij}]_{i,j=1}^{n}$ has a non-zero null vector $v$. Consider the relation
$(1-\overline{z}_{i}z_{j})\widetilde{g}_{ij}=(1-r^{2})\widetilde{g}_{ij}+(r^{2}-\overline{z}_{i}z_{j})\widetilde{g}_{ij}.$
Since $\widetilde{g}_{ij}>0$, $(1-r^{2})\widetilde{g}_{ij}>0$ and, by equation (6.11), $(r^{2}-\overline{z}_{i}z_{j})\widetilde{g}_{ij}\geq 0$, which is a contradiction. ∎
By Theorem 5.12, if a DP Pick problem is solvable, then there exists a rational solution $\varphi\in\mathcal{S}_{\mathrm{dp}}$. In the next theorem we show that if, further, the problem is extremally solvable then there exists $T\in\mathcal{F}_{\rm dp}(\delta,\lambda)$, acting on an $n$-dimensional Hilbert space, such that $\|\varphi(T)\|=\|\varphi\|_{\rm dp}=1.$
@@
Let us now prove the existence of an operator $T$ with the stated properties. By assumption, the DP-Pick problem $\lambda_{j}\mapsto z_{j},j=1,\dots,n$ is extremally solvable. By Theorem 6.3, there exists $\widetilde{g}\in\mathcal{G}_{\rm dp}(\lambda)$ such that
$\operatorname{rank}\big[(1-\overline{z}_{i}z_{j})\widetilde{g}_{ij}\big]_{i,j=1}^{n}<n,$ (6.16)
so that $\begin{bmatrix}(1-\overline{z}_{i}z_{j})\widetilde{g}_{ij}\end{bmatrix}$ is singular, and therefore has a non-zero null vector $\xi=[\xi_{1},\dots,\xi_{n}]^{T}\in\mathbb{C}^{n}$, which is to say that
$\sum_{j=1}^{n}(1-\overline{z}_{i}z_{j})\widetilde{g}_{ij}\xi_{j}=0\ \text{for}\ i=1,\dots,n.$ (6.17)
Since $\widetilde{g}\in\mathcal{G}_{\rm dp}(\lambda)$, $\widetilde{g}_{ij}]>0$, and so $\widetilde{g}_{ij}]$ has rank $n$. By Moore’s theorem’s Theorem there exist an $n$-dimensional Hilbert space $\mathcal{H}$ and a basis $\widetilde{e}_{1},\dots,\widetilde{e}_{n}\in\mathcal{H}$ such that $\widetilde{g}_{ij}=\langle\widetilde{e}_{j},\widetilde{e}_{i}\rangle$ for $i,j=1,\dots,n$.
Define an operator $T$ on $\mathcal{H}$ by $T\widetilde{e}_{j}=\lambda_{j}\widetilde{e}_{j}$ for $j=1,\dots,n$. Since $\widetilde{g}\in\mathcal{G}_{\rm dp}(\lambda)$, by Proposition 4.10, $T\in\mathcal{F}_{\rm dp}(\delta,\lambda)$. Note that $\varphi(T)\widetilde{e}_{j}=z_{j}\widetilde{e}_{j}$, and so, if $x=\sum_{j=1}^{n}\xi_{j}\widetilde{e}_{j}$, then
$\varphi(T)x=\sum_{j=1}^{n}z_{j}\xi_{j}\widetilde{e}_{j}$
@@
$\langle(1-\varphi(T)^{*}\varphi(T))x,x\rangle$ $=\sum_{i,j=1}^{n}(1-\overline{z}_{i}z_{j})\overline{\xi}_{i}\xi_{j}\langle\widetilde{e}_{j},\widetilde{e}_{i}\rangle$
$=\sum_{i,j=1}^{n}(1-\overline{z}_{i}z_{j})\overline{\xi}_{i}\xi_{j}\widetilde{g}_{ij}$
$=\sum_{i=1}^{n}\overline{\xi}_{i}\sum_{j=1}^{n}(1-\overline{z}_{i}z_{j})\xi_{j}\widetilde{g}_{ij}$
@@
As $\xi\neq 0$, the complex numbers $\xi_{1},\dots,\xi_{n}$ are not all zero, and so, since $\widetilde{e}_{1},\dots,\widetilde{e}_{n}$ are linearly independent, $x=\sum_{j=1}^{n}\xi_{j}\widetilde{e}_{j}\neq 0$. Since $\|\varphi\|_{\mathrm{dp}}\leq 1$ and $T\in\mathcal{F}_{\rm dp}(\delta,\lambda)$, we have $\|\varphi(T)\|\leq 1$, and so $1-\varphi(T)^{*}\varphi(T)\geq 0$. In conjunction with the equality (6), this implies that $(1-\varphi(T)^{*}\varphi(T))x=0$, and hence $\|\varphi(T)x\|^{2}=\|x\|^{2}$. Since $\neq 0, xx$ is a maximizing vector for $\varphi(T)$ and $\|\varphi(T)\|=1$.
@@
where $\dim\mathcal{M}\leq 2n$. Since $\in\mathcal{F}_{\rm dp}(\delta,\lambda), TT$ satisfies
@@
$1-\varphi(T)^{*}\varphi(T)=u(T)^{*}\big(1-E(T)^{*}E(T)\big)\ u(T).\qed$
7. Declarations
NSF and EPSRC grants. Agler was partially supported by National Science Foundation Grants DMS 1361720 and 1665260. Lykova and Young were partially supported by the Engineering and Physical Sciences grant EP/N03242X/1.
Conflict of interest. The authors have no Conflict of interest to declare that are relevant to the content of this article.
Data availability statement. No data were collected, generated or consulted in connection with this research.